OpenStudy (anonymous):

Simply the expression. Write your answer in lowest terms and leave it in factored form. 5y/ y(2y-1) + 2 / 2y-1

OpenStudy (mertsj):

What is the common denominator?

OpenStudy (mertsj):

\[\frac{5y}{y(2y-1)}+\frac{2}{2y-1}\]

OpenStudy (anonymous):

the common denominator is 2y-1

OpenStudy (mertsj):

For the first denominator, we need the factors y and 2y-1. For the second denominator we need 2y-1 but we already have it from the first denominator. So the common denominator is y(2y-1).

OpenStudy (mertsj):

So write the second fraction with the common denominator and leave the first fraction alone since it already has the correct denominator.

OpenStudy (anonymous):

2/ y(2y-1)

OpenStudy (mertsj):

\[\frac{2}{2y-1}\times\frac{y}{y}=\frac{2y}{y(2y-1)}\]

OpenStudy (anonymous):

okay, where does that leave the 5y?

OpenStudy (mertsj):

\[\frac{5y}{y(2y-1)}+\frac{2y}{y(2y-1)}=\frac{7y}{y(2y-1)}\]

OpenStudy (anonymous):

okay. thank you so much!

OpenStudy (anonymous):

Hope I can have more of your help!

OpenStudy (mertsj):

Now reduce the final answer by cancelling out the common factor y

OpenStudy (anonymous):

7/ (2y-1)

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

Finally, Something clicked! Now to find more help for my other problems.

OpenStudy (mertsj):

You might surprise yourself and not need help.

OpenStudy (anonymous):

Oh no, every time something "clicks" I end up failing my tests.

OpenStudy (mertsj):

Practice, practice, practice